OpenStudy (anonymous):

k-8=7 i would say 8+7 then

OpenStudy (anonymous):

8-7

OpenStudy (anonymous):

Yes, you're right k = 8+7, you see what you do to one side you must also do the other, the goal is to get the variable on its own meaning by itself, in this case k is your variable and you want it alone. Generally, in algebra, all you have to do is use the opposite operation of which already exists, so for example yours says k-8=7. You have a -8, so what would eliminate it? +8 right? Lets try: \[k-8+\color{red}{8}=7+\color{red}{8} \implies k =7+8 \implies 15\]

OpenStudy (anonymous):

You can see the 8s get cancelled on the left side and you're only left with it on the right.